kinh:

What is the factorization of 729x^15 + 1000? *question has been edited by a moderator for clarity *

Vocaloid:

if you rewrite 729, x^15, and 1000 as a product of perfect cubes you have (9x^5)^3 + 10^3 now you have a perfect cube expression a^3 + b^3 where a = (9x^5) and b = 10 now you can factor it out as a^3 + b^3 = (a+b)(a^2- 2ab + b^2) just replace a and b with the appropriate values lmk if you are stuck or want to check your solution welcome to QC~